# Prague, Czech

### 10 Mar 2026 v1

@@ 5 Limitations
Our approach relies on the quality of the training data: the closer the distributions of the training and test sets, the better the model’s accuracy. Ideally, the test data should follow a distribution similar to that used to train the model; however, this is not strictly required, as we have not yet evaluated the model’s generalization ability outside this scenario. Future studies will investigate which factors (e.g., matrix sizes or matrix element magnitude) are most sensitive and contribute to performance loss on testing data. In this paper, we focus on the formulation of our approach and demonstrate the capabilities of GNNs for condition number estimation. We have not optimized the model architecture, hyperparameters, or training routine, so readers may achieve further refined results by tuning these settings.
@@ 6 Conclusion
We proposed a rapid condition number estimation approach using graph neural networks for sparse linear systems. We propose two prediction schemes, one by decomposing the condition number into an exactly computable forward norm and a neural-network-predicted inverse norm, and another by directly predicting the matrix condition number. Extensive experiments for the two schemes show that our method achieves orders of magnitude speedup over exact computation and compares favorably against the Hager-Higham estimator at a certain scale in terms of speed while maintaining competitive accuracy. The architecture combines graph convolutional layers to capture sparsity patterns with global feature encoders for statistical properties, enabling effective generalization. Future work will explore optimized GNN architectures and end-to-end feature engineering.
We believe that our work can be beneficial to existing work on precision tuning (e.g. Carson and Chen (2026)), since precision usage is highly dependent on condition number and a faster estimation of it can significantly boost the applications.

### 15 Mar 2026 v2

@@ 5 Limitations
Our approach relies on the quality of the training data: the closer the distributions of the training and test sets, the better the accuracy of the model. Ideally, the test data is expected to follow a distribution similar to that used to train the model; however, this is not strictly required, as we have not yet evaluated the model’s generalization ability outside this scenario. Future studies will investigate which factors (e.g., matrix sizes or matrix element magnitude) are most sensitive and contribute to performance loss on testing data. In this paper, we focus on the formulation of our approach and demonstrate the capabilities of GNNs for condition number estimation. We have not optimized the model architecture, hyperparameters, or training routine, so readers may achieve further refined results by tuning these settings. Besides, as observed in the experiments, the dominant runtime for condition number prediction for GNN lies in the inference. Therefore, potential speedup can be achieved with ONNX [3].
@@ 6 Conclusion
We proposed a rapid condition number estimation approach using graph neural networks for sparse linear systems. We propose two prediction schemes, one by decomposing the condition number into an exactly computable forward norm and a neural-network-predicted inverse norm, and another by directly predicting the matrix condition number. Extensive experiments for the two schemes show that, with a limited amount of data, our method achieves orders of magnitude speedup over exact computation and compares favorably against the Hager-Higham and Golub–Kahan methods at a certain scale in terms of speed while maintaining competitive accuracy. The architecture combines graph convolutional layers to capture sparsity patterns with global feature encoders for statistical properties, enabling effective generalization. Future work will explore optimized GNN architectures and end-to-end feature engineering.
We believe that our work can be beneficial to existing work on precision tuning (e.g. [1]), since precision usage is highly dependent on condition number and a faster estimation of it can significantly boost the applications.
Appendix A Loss during training
(a) 1-norm ($g(\mathbf{A})\approx\|\mathbf{A}^{-1}\|$)
(b) 2-norm ($g(\mathbf{A})\approx\|\mathbf{A}^{-1}\|$)
(c) 1-norm ($g(\mathbf{A})\approx\kappa(\mathbf{A})$)
(d) 2-norm ($g(\mathbf{A})\approx\kappa(\mathbf{A})$)
Figure 4: Loss during training.